OpenStudy (anonymous):

What is the LCM of 15 and 60?

geerky42 (geerky42):

HINT: \(15\times4 = 60\)

OpenStudy (yanasidlinskiy):

Lowest common factor as I've showed you in the other post:)

OpenStudy (anonymous):

3.. but really??? 15 divided by 3 is 5 60 divided 3 is 20

OpenStudy (shadowlegendx):

The Least Common Multiple is the lowest multiple of both numbers that isn't zero or itself.

OpenStudy (shadowlegendx):

In this case anyway(the both numbers part).

OpenStudy (mathstudent55):

The least common multiple of two numbers, one of which is a multiple of the other, is the larger of the two numbers.

OpenStudy (mathstudent55):

The GCF is 15, not 3.

OpenStudy (anonymous):

60. It is both 60 and 15 least common multiple

geerky42 (geerky42):

Figures, @Rapunzel101402 only wants answer...

OpenStudy (anonymous):

stop

OpenStudy (skullpatrol):

$$\Huge60 = 2^2*3*5$$$$\Huge15=3*5$$$$\Huge\text{LCM =} 2^2*3*5=60$$